OpenStudy (anonymous):

9-M OVER 3=2 HELP PLEASE

OpenStudy (anonymous):

\[(9 - M) \over 3\] = 7

OpenStudy (anonymous):

multiply both sides by 3 \[3[(9-M)/3] = 3[7]\]

OpenStudy (anonymous):

9 - M = 21

OpenStudy (anonymous):

-M = 12 M = -12

OpenStudy (anonymous):

WAIT WAT

OpenStudy (anonymous):

what don't you understand?

OpenStudy (anonymous):

sorry...typed that into wrong question :P

OpenStudy (anonymous):

what do you need chocolate?

OpenStudy (anonymous):

oh wait, now I see... I confused 7 with 2

OpenStudy (anonymous):

hold on a sec, I will fix the equation

OpenStudy (anonymous):

\[3[(9-M)/3)] = 3(2)\]

OpenStudy (anonymous):

giving you 9 - m = 6

OpenStudy (anonymous):

-M = -3 and M = 3

OpenStudy (anonymous):

final answer.

OpenStudy (anonymous):

sorry for confusion

OpenStudy (anonymous):

thats ok